Solve the problem and show the complete solution.

A uniform electric field E = 5000 N/C is passing through a flat square area A= 2m2 . Determine the electric flux. θ = 60o (the angle between the electric field direction and a line drawn perpendicular to the area)


1
Expert's answer
2021-02-15T00:55:27-0500

The electric flux of the uniform field is given as follows:


"\Phi = E_nA"

where "E_n" is the normal component of the electric field, and "A = 2m^2" is the area. Since the angle between the electric field direction and a line drawn perpendicular to the area is "\theta = 60\degree", the normal component of the field "E = 5000N\/C" is:


"E_n = E\cos\theta"

Finally, obtain:


"\Phi = AE\cos\theta\\
\Phi = 2\cdot 5000\cdot \cos60\degree = 5000\space \dfrac{N\cdot m^2}{C}"

Answer. "5000\space \dfrac{N\cdot m^2}{C}".